OpenStudy (hunyx3):

The scale of a map is 1 cm : 73 km. What is the actual distance between two towns that are 4 cm apart on the map? The distance between he cities is ___ km.

jhonyy9 (jhonyy9):

use the rule of 3 simple 1 cm ---- 73 km 4 cm ---- x km ----------------- x = 4*73/1 = 4*73 = ? km